OpenStudy (anonymous):

Ok I just need to know how to set this problem up and then I can solve it. One canned juice drink is 15% orange juice; another is 5% orange juice. How many litres of each should be mixed together in order to get 10L that is 7 % orange juice. I have to use a system of equations to solve this. Does x+y =7 and then 15x + 5y = (10)(7) make sense? Somebody please help I am still stuck on this question I have to use a system of equations to answer this. The equations I am using are not giving me the right answer.

OpenStudy (anonymous):

\[.15x+.05(10-x)=.07\times 10\]

OpenStudy (anonymous):

x+y=10 5x+15y=70

OpenStudy (anonymous):

you do not need a "system"

OpenStudy (anonymous):

yes you do

OpenStudy (anonymous):

no, you do not

OpenStudy (anonymous):

elixated I tried that system and it gave me a negative answer and that can't be right

OpenStudy (anonymous):

you can solve it without a system but to get the quanity x-10 you should start witha system

OpenStudy (anonymous):

I give u equation early percent ---------------- quantity Juice I 15 ---------------- x liters JuiceII 5 ---------------- 10 - x liters Total 7.00% ---------------- 10 15x+5(10-x)=7*10 15x+50-5x=70 15x-5x=70-50 10x=20 /10 x=2 liters 15.00% 8liters 5.00%

OpenStudy (anonymous):

i just solved it without one. put the amount of 15% juice as say X. then the amount of 5% juice is 10-x

OpenStudy (anonymous):

You did something wrong. When you solve it you should get a positive answer. I did.

OpenStudy (anonymous):

ok here we go

OpenStudy (anonymous):

x + y = 10 0.15x + 0.05y = (10)(0.07) 0.05y = -0.15x + 0.7 Y = -3x + 0.7 X – 3x +0.7 = 10 -2x + 0.7 = 10 -2x = 9.3 I didn't finish solving because I knew the answer would be negative

OpenStudy (anonymous):

\[.15x+.05(10-x)=.07\times 10\] \[15x+5(10-x)=70\] \[15x+50-5x=70\] \[10x+50=70\] \[10x=70-50=20\] \[x=\frac{20}{10}=2\]

OpenStudy (anonymous):

Ok thank you that helps a lot

OpenStudy (anonymous):

so i get 2 liters of 15% juice and therefore 10-2=8 liters of 5% juice. lets check the answer and see if it is right

OpenStudy (anonymous):

you guys cannot use decimals. instead, set up your equation as: x+y=10 5x+15y=70

OpenStudy (anonymous):

\[2\times .15+8\times .05=.3+.4=.7\] and .7 is 70% of 10 so that is the answer

OpenStudy (anonymous):

15y = -5x + 70 y = -1/3x + 70 x -1/3x + 70 = 10 2/3x + 70 = 10 2/3x = -60 answer will still be negative

OpenStudy (anonymous):

"you guys cannot use decimals"??? can we use fractions? how about variables?

OpenStudy (anonymous):

where on earth did \[15y=-5x+70\] and \[y=-\frac{1}{3}x+70\] come from?

OpenStudy (anonymous):

because 5x + 15y = 70 -5x -5x 15y = -5x + 70 /15 /15 forgot to divide 70 by 15 y = -1/3x + 4 2/3

OpenStudy (anonymous):

or 14/3

OpenStudy (anonymous):

but then it would be X + -1/3x + 14/3 = 10 2/3x + 14/3 = 10 2/3x + 14/3 = 30/3 2/3x = 16/3 x = 48/6 = 8 so not negative sorry